# Constructing Benchmarks and Interventions for Combating Hallucinations in LLMs

### v1

@@ 4.2 Evaluating Intervention Impact
Recent work lacks a consistent evaluation method for measuring intervention success. Most compare probabilities of the golden answer and other possible answers (Marks & Tegmark, 2023; Zou et al., 2023; Li et al., 2023; Zhang et al., 2024; Chen et al., 2024; Hoscilowicz et al., 2024), while others also generate answers and evaluate them with a language model (Li et al., 2023; Zhang et al., 2024; Chen et al., 2024).
We advocate for examining multiple metrics to provide a holistic picture of intervention effects. To measure intervention success, we report the classification accuracy. In the open-book case, we measure how often the model prefers the contextual answer over the parametric answer by comparing Rank()Rank(). In the closed-book case, we compare the rank of the golden answer given a bad-shot prompt and a good-shot prompt (Rank bad shotsRank good shots). See Appendix C for detailed information.
The classification accuracy is simple and intuitive. However, recent work suggests that comparing probabilities of answer tokens is not a good proxy for text generation (Wang et al., 2024). We argue that this discrepancy is particularly crucial for intervention assessment. Therefore, we report the generation accuracy, which examines whether the desired answer was generated in greedy decoding of 20 tokens. Note, that we focus on a generation of a few tokens, as the model may produce a lengthy correct answer that should also be classified as grounded.
Finally, to ensure that the intervention has not compromised the model’s performance, we report its perplexity on 100 random articles from WikiText-103 (Merity et al., 2016) before and after an intervention. We call this Wiki ppl. See additional details in Appendix C.777We also tried pure perplexity but it was less stable than Wiki ppl.
@@ 5 Implementation Details
In all experiments, we randomly selected 1000 examples from each label (hallucinate/grounded) for analysis in each dataset, and split to 70%/10%/20% for training/validation/test.888Result on the general distribution of the data show similar but weaker trends; see Appendix F.4. We use a linear classifier (SVM) for detection, similar to prior work (Li et al., 2023; Marks & Tegmark, 2023). All results are averaged across three experiments, each with a different random seed, random split into training, validation, and testing, and the classifier seed in detection experiments. Detection results and the generalization of the detection in different settings can be found in Appendices D and F.5.
We evaluate Llama-2-7B (Touvron et al., 2023) in most experiments. We also evaluate GOAT-7B (Bekbayev et al., 2023), a fine-tuned version of Llama-2, in a subset of the experiments, to highlight differences between base and fine-tuned models.
To explore the various options raised in Section 4.1, we start with mass-mean direction intervention with dynamically and pre-answer input to , a configuration that showed promising results in initial exploration. We then modify a single variable at a time (dynamic/static intervention, computing the steering vector pre/post the answer etc.) to assess its impact on performance in an ablation setup. Additional hyper-parameter details are in Appendix E.999When comparing interventions with mass-mean vs. classifier direction, we found inconsistent results with no clear pattern. See discussion in Appendix F.3..
@@ 6 Intervention Results
Figures 6(a) and 7(a) show the accuracy on the open-book DisentQA and closed-book TriviaQA settings w.r.t intervening with increasing values of , which controls the magnitude of the intervention. We see different behaviors in the two settings. While in closed-book classification and generation accuracy are coupled, in open-book, when is increased, classification accuracy mostly increases but generation peaks and then drops, indicating that classification accuracy is not always a good approximation for generation accuracy, which is what we care for when using the model language abilities. We also see drop in Wikipedia perplexity as grows (Figures 6(c) and 7(c)), indicating worse language modeling capability with stronger intervention. This occurs in both settings but with different changes per setting. Therefore, evaluating also perplexity is recommended when assessing intervention performance.
@@
Component-wise, we see that closed-book and open-book also behave differently. In open-book, attention and heads are best performing and very similar, under classification, generation and perplexity evaluation. In closed-book, residual has better classification and generation accuracy, but it declines at higher and perplexity levels are increasing quickly too, while attention accuracy keeps improving without any perplexity drop.
@@
In Section 4.1 we proposed a novel pre-answer intervention strategy. Figure 8 presents a comparison between pre-answer and post-answer intervetion strategies101010We run this test only on the open-book setting as it has a clear answer ( for grounded examples and for hallucinate examples).. We can see that while on classification accuracy, the two strategies are on par, under generation and perplexity, it is clear that pre-answer is a better strategy. In addition, under post-answer, intervention at the heads component is best performing, but under pre-answer, attention component gains and matches heads performance.
@@
We conclude that a fine intervening component in both settings is the attention, and to reach good results intervention should be performed using the pre-answer strategy.
6.1 Static Interventions vs. Dynamic Intervention
@@
This section compares static and dynamic interventions. As discussed in Section 4.1, static intervention is the common approach in the literature and modifies the same components in every input example. This intervention usually selects the places with the highest detection to always intervene in. Dynamic intervention is a new approach we propose, which decides which components to modify given an input example based on the detector’s decision. Dynamic intervention might be preferable since unnecessary interventions might do more harm than good (the steering vector is designed to move the representation between subspaces, using it is not designed for cases where the representation is already in the "correct" subspace).
@@
As Table 1 shows, in both settings, there is minimal difference between the two approaches in most cases with a slight overall preference for the static intervention. However, when intervening in the residual, dynamic intervention yields significantly better results, while static intervention leads to poor generation accuracy and perplexity. Again, we see that classification accuracy is not correlated with generation accuracy, this time in both settings.
To further illustrate the effect on the residual component, we provide examples of generated text using static vs. dynamic interventions that demonstrate how static intervention in the residual component negatively impacts generation in Appendix F.1.
We conclude that when intervening in the residual it is preferred to use the dynamic intervention and elsewhere it varies with generally slight preference for the static intervention.
| Setting | DisentQA | TriviaQA | |||||
| component | direction type | Classification Accuracy | Generation Accuracy | Wikipedia Perplexity | Classification Accuracy | Generation Accuracy | Wikipedia Perplexity |
@@ 6.2 Finetuning Facilitates Intervention
In this section, we explore the impact of fine-tuning on intervention success by comparing two models: Llama2-7B vs. Goat-7B, a fine-tuned version of Llama2.111111We decided to leave chat models for future work as they require modifying the prompt to a user/assistant format instead of a regular prompt, making direct comparisons of base and chat models difficult. Table 2 shows the average change resulting from intervention compared to no intervention. (Full results are found in Appendix F.2.) Notably, the intervention significantly enhances the generation and classification accuracy of Goat compared to Llama2 across most components, without negatively affecting perplexity. This suggests the advantage of utilizing the fine-tuned model over its basic counterpart, as shown in Li et al. (2023) but does not correspond to other works (Zhang et al., 2024). More research is needed to characterize differences in intervention effects on pre-trained and fine-tuned models.
| Setting | DisentQA | TriviaQA | |||||
| component | direction type | Classification Accuracy | Generation Accuracy | Wikipedia Perplexity | Classification Accuracy | Generation Accuracy | Wikipedia Perplexity |
@@ 7 Related Work
Prior work has considered cases of LM failures under various terms such as factuality (Chuang et al., 2023; He et al., 2023), lying (Azaria & Mitchell, 2023; Pacchiardi et al., 2023), truthfulness (Marks & Tegmark, 2023; Li et al., 2023; Zhang et al., 2024; Zou et al., 2023), sycophancy (Sharma et al., 2023), and as hallucinations (Shi et al., 2023; Dhuliawala et al., 2023). Hallucinations are sometimes defined as cases of model mistakes that seem plausible to a user (Ji et al., 2023).
An important distinction is between extrinsic (closed-book) and intrinsic (open-book) hallucinations (Ji et al., 2023; CH-Wang et al., 2023). The interventionist literature has mostly studied the closed-book case, while we consider both cases.
Research on hallucination typically falls into two main categories: detection and mitigation. Detecting hallucinations can involve treating the model as a black box, and posing questions or sampling (Gekhman et al., 2023; Pacchiardi et al., 2023; Manakul et al., 2023). Another line of work has tried to detect hallucinations or factuality by examining the model’s hidden representations, often by training a detection classifier (Burns et al., 2022; He et al., 2023; Rateike et al., 2023; Slobodkin et al., 2023; Azaria & Mitchell, 2023; CH-Wang et al., 2023; Yuksekgonul et al., 2023; Chen et al., 2023; Yin et al., 2024) We similarly use detectors in this work, with an emphasis on detection before a hallucination occurs.
Mitigating hallucinations can be done via prompting (Si et al., 2022; Dhuliawala et al., 2023; Lei et al., 2023) or fine-tuning (Tian et al., 2023; Razumovskaia et al., 2023). Other methods involve modifying the model’s logits (Chuang et al., 2023; Kalai & Vempala, 2023). Our work focuses on addressing hallucinations by intervening in the inner layers of the model (Zhang et al., 2024; Li et al., 2023; Marks & Tegmark, 2023; Zou et al., 2023; Chen et al., 2024; Hoscilowicz et al., 2024), which potentially does not require external resources or longer prompts. In this work, we characterize the effect of different choices for intervention, both from prior work and novel techniques we propose.
@@ 8 Discussion and Conclusion
In this work, we established a framework for assessing white-box hallucination mitigation techniques in open-book and closed-book settings. We propose a typology of hallucination types from the lens of a model’s knowledge, and highlight the feasibility of type-3 hallucinations – cases where the model knows the answer but does not generate it. We discussed the significance of these aspects and developed a methodology for creating appropriate datasets with hallucination labels.
We performed a comprehensive investigation of different variables involved in interventionist approaches for mitigating hallucinations. Notable insights include the importance of computing intervention vectors before reading an answer (input to ); the difference between evaluating classification and generation accuracy and the importance of perplexity evaluation; the potential of dynamic interventions that vary by example with its main importance in residual intervention; and the pros and cons traits of different components – while intervening in the residual reduces hallucinations, it also compromises the model’s language modeling capabilities. In comparison, intervention in the attention component consistently performs well across various measures and datasets.
This study establishes a framework for evaluating white-box hallucination mitigation techniques across open-book and closed-book settings, including a typology of hallucination types based on the model’s knowledge, while uncovering insights into effective intervention strategies.
9 Limitations
This work considers a wide range of possible configurations when mitigating hallucinations via white-box interventions. However, one limitation is that we only focused on one model (Llama2-7B) and one of its fine-tuned variants (Goat-7B). Additionally, we presented results on one open-book dataset (DisentQA) and one closed-book dataset (TriviaQA), both in question-answering setups. Lastly, we use an initial setting and modify a single variable each time instead of checking across all options of modifications. To form a broader understanding it is essential to investigate additional settings and models.
@@ A.1 General Dataset Construction Specifics
We created at least 1,000 examples labeled as “hallucinate" and 1,000 as “grounded". We filtered out instances where the golden/parametric/contextual answers exceeded 5 tokens or where the prompt exceeded 1,000 tokens.
To assess the model’s knowledge of the golden answer, we employed a beam search generation of 20 tokens, repeated 5 times with a temperature of 0.8. We utilized only examples falling under type 3 (Section 3.1), ensuring each answer was generated at least four times.
@@ A.3 Closed-Book Dataset Construction Specifics
| Prompt | Grounded | Hallucinated |
@@ Appendix B Final Datasets Statistic
In section 3 we created a labeled dataset given a model and a dataset. DisentQA for the open-book dataset and TriviaQA for the closed-book dataset. The statistic of the final dataset count of the number of examples in each label can be found in Tabel 4, the else label is all the examples that did not classify to the other two labels. We can see that the different settings have different results, indicating the importance of creating a labeled dataset regarding the specific model.
| Data name | # Grounded | # hallucinate | # else |
| closed book-Llama2-TriviaQA | 4553 | 1451 | 10306 |
| open book-Llama2-DisentQA | 1673 | 2869 | 8958 |
| closed book-Goat-TriviaQA | 2618 | 1507 | 12185 |
| open book-Goat-DisentQA | 3019 | 1261 | 9220 |
@@ Appendix C Specific Evaluation Measures
In the open book setting, we assess classification accuracy by comparing the model’s preference for the , Rank()Rank() Regarding the closed book setting, we compare Rank( | bad shots) Rank( | good shots). In both settings, the rank calculation mirrors that of the dataset creation step. If the equations are true the example is labeled ‘grounded’, else ‘hallucinate’.
@@
Lastly, to calculate the Wiki ppl we use 100 random articles and from each, we use the first 100 tokens to calculate the perplexity.
@@ Appendix D Detection Results
Figures 9(a) and 9(b) show hallucination detection results on the (open-book) DisentQA-based dataset and (closed-book) TriviaQA-based dataset, respectively, when applied to different model components at different layers, using Llama-2-7B.
On most layers, detection is significantly higher than random (dashed lines), indicating a signal of possible hallucination prior to the occurrence of hallucination. Additionally, we observe that the different components yield fairly similar results, with a slight advantage for detection based on the residual stream.
We remind the reader that detecting possible hallucinations before they occur is a more challenging task than detecting them post-hoc (CH-Wang et al., 2023). Indeed, if we concatenate the answer to the prompt, we reach over 90% accuracy in most layers and components, see Figure 13. Given that in practical use cases the answer is not known ahead of generation, we emphasize the importance of pre-hoc detection.
For completeness, we present the detection results on one data split of the validation set on the residual, attention, and MLP layers in Figures 10 and 11, utilizing three different classifiers: a linear classifier trained on the vectors as they are (labeled classifier in the figures), a linear classifier trained only on the vectors’ norms (labeled norm in the figures), and a linear classifier trained on the normalized vectors (labeled direction in the figures). It’s evident that the linear classifier trained on the normalized vectors (direction classifier) achieves the best results across most components, layers, and datasets.
Additionally in Figure 12 we see the classification results on the heads on Llama2-7B in the two settings. We can see that those results are a bit less good than the ones on the other model components with much better results in the open-book setting than the closed-book setting.

### v2

@@ 4.2 Evaluating intervention impact
Recent work lacks a consistent evaluation method for measuring intervention success. Most compare probabilities of the golden answer and other possible answers [24, 51, 22, 50, 7, 12], while others also evaluate generation with a language model [22, 50, 7]. We advocate for examining multiple metrics to provide a holistic picture of intervention effects.
To measure intervention success, we report the classification accuracy. In the open-book case, we measure how often the model prefers the contextual answer over the parametric answer by comparing Rank()Rank(). In the closed-book case, we compare the rank of the golden answer given a bad-shot prompt and a good-shot prompt (Rank bad shotsRank good shots).
The classification accuracy is simple and intuitive. However, recent work suggests that comparing probabilities of answer tokens is not a good proxy for text generation [41]. We argue that this discrepancy is particularly crucial for intervention assessment. Therefore, we report the generation accuracy, which examines whether the desired answer was generated in greedy decoding of 20 tokens. Note, that we focus on a generation of a few tokens, as the model may produce a lengthy correct answer that should also be classified as grounded.
@@ 5 Implementation details
In all experiments, we randomly selected 1000 examples from each label (hallucination/grounded) for analysis in each dataset and split to 70%/10%/20% for training/validation/test. We use a linear SVM classifier for detection, as in prior work [22, 24]. Each experiment was repeated with three random seeds for the SVM and split into training/validation/test sets. We report average results with standard deviations. Datasets construction and all the results ran on NVIDIA RTX 6000 Ada (49GB) with 4 CPUs. Each dataset construction and each result takes approximately 12 hours (30 GPU days).
In addition to evaluating Llama-2-7B [39] on all experiments, we evaluated Goat-7B [3] on a subset of the experiments to highlight performance differences between base and fine-tuned models.888We leave chat models for future work as they modify the prompt to a chat format that makes direct comparisons of base and chat models difficult.
To explore the various options discussed in Section 4.1, we start with dynamic pre-answer mass-mean direction intervention with , a configuration that showed promising results in initial exploration. We then modify a single intervention variable at a time (dynamic/static, pre-answer/post-answer, etc.) to assess its impact on performance in an ablation setup. Additional hyperparameter details, ablation choices, and results on the general distribution of the data are in Appendices E, F.3 and F.4.
@@ 6 Intervention results
Figures 6(a), 6(b), 7(a) and 7(b) show the accuracy on the open-book DisentQA-WACK-Llama2 (DisentQA-WACK for short) and closed-book TriviaQA-WACK-Llama2 (TriviaQA-WACK for short) settings w.r.t intervening with increasing values of , which controls the magnitude of the intervention. While in closed-book setting, classification and generation accuracy are coupled, in open-book, when is increased, classification accuracy mostly increases but generation peaks and then drops. This indicates that classification accuracy is not always a good approximation for generation accuracy, which is what a user actually receives from an LLM. We also see an increase in Wiki ppl as grows (Figures 6(c) and 7(c)), pointing at worse language modeling ability with stronger intervention. This occurs in both settings but with different impacts per intervened component. Thus, evaluating also perplexity is recommended to assess intervention performance.
@@
Component-wise, in open-book, attention and heads are best performing and very similar, under classification, generation, and perplexity evaluation. On the other hand, in closed-book, residual has better classification and generation accuracy, but declining at higher , and perplexity levels are increasing quickly too, while attention accuracy keeps improving without any perplexity drop.
@@
In Section 4.1 we proposed a novel pre-answer intervention strategy. Figure 8 presents a comparison between pre-answer and post-answer interventions.999We run it only on the open-book setting as it has a clear answer ( for grounded and for hallucination). We can see that while the two strategies are on par on classification accuracy, regarding generation, and perplexity, it is clear that pre-answer is a better strategy. In addition, under post-answer, intervention at the heads component is best performing, but under pre-answer, attention component gains and matches heads performance.
@@
We conclude that a recommended intervening component in both settings is the attention, and to reach good results intervention should be performed using the pre-answer strategy.
6.1 Static interventions vs. dynamic interventions
@@
Static intervention, which modifies the same components for every input and typically targets areas with the highest detection, is the common approach in the literature. Dynamic intervention, which we propose, decides which components to modify given an input example based on the detector’s decision. Dynamic intervention might be preferable since unnecessary interventions might do more harm than good, e.g., hurting generation abilities without improvement to “correctness”.
@@
Table 1 shows a minimal difference between the two approaches in most cases with a slight overall preference for the static intervention. However, when intervening in the residual, dynamic intervention yields significantly better results, while static intervention leads to poor generation accuracy and perplexity. Additionally, examples of generated text (in Appendix F.1) illustrate how static intervention in the residual component adversely affects generation. We conclude that when intervening in the residual, it is preferred to use the dynamic intervention.
| Setting | DisentQA-WACK | TriviaQA-WACK | |||||
| Classification | Generation | Classification | Generation | ||||
| Accuracy | Accuracy | Wiki ppl | Accuracy | Accuracy | Wiki ppl | ||
@@ 6.2 Finetuning facilitates intervention
We explore the impact of fine-tuning on intervention success by comparing Llama2-7B vs. Goat-7B, a fine-tuned version of Llama2. Table 2 shows the average change resulting from intervention compared to no intervention (Full results are in Appendix F.2). The intervention significantly enhances Goat’s generation and classification accuracy compared to Llama2 across most components, without worsening perplexity. This suggests the advantage of utilizing the fine-tuned model over its basic counterpart, as shown in [22] but is not consistent across all studies [50] and requires further research.
| Setting | DisentQA-WACK | TriviaQA-WACK | |||||
| Classification | Generation | Classification | Generation | ||||
| Accuracy | Accuracy | Wiki ppl | Accuracy | Accuracy | Wiki ppl | ||
@@ 7 Related work
Prior work has considered cases of LM failures under various terms such as factuality [8, 11], lying [1, 29], truthfulness [24, 22, 50, 51], sycophancy [33], and hallucinations [34, 9]. Hallucinations are sometimes defined as cases of model mistakes that seem plausible to a user [15].
Research on hallucination typically falls into two main categories: detection and mitigation. Detecting hallucinations can involve treating the model as a black box, and posing questions or sampling [10, 29, 23]. Another line of work has tried to detect hallucinations or factuality by examining the model’s hidden representations, often by training a detection classifier [4, 11, 30, 36, 1, 5, 48, 6, 46, 21] We similarly use detectors, emphasizing detection before a hallucination occurs.
Mitigating hallucinations can be done via prompting [35, 9, 20] or fine-tuning [38, 32] or involve modifying the model’s logits [8, 19]. Our work focuses on addressing hallucinations by intervening in the inner layers of the model [50, 22, 24, 51, 7, 12], which potentially does not require external resources. One difference lies in closed-book versus open-book hallucinations [15, 5]. Intervention body of work mostly studied the closed-book case, while we consider both cases, and characterize the effect of different choices for intervention, both from prior work and novel techniques we propose.
@@ 8 Discussion and conclusion
This work presented WACK, a framework for assessing white-box hallucination mitigation techniques in open-book and closed-book settings. We propose a typology of hallucinations through the lens of a model’s knowledge. We highlight the possibility of hallucinations even though the model knows the answer, and propose an automatic way of creating a labeled dataset of such hallucinations.
We then performed a comprehensive investigation of different variables involved in interventionist approaches for mitigating hallucinations. Notable insights include the importance of computing intervention vectors using pre-answer vectors instead of post-answer vectors; showing that the novel dynamic intervention approach allows robust interventions; attention intervention consistently performs well across various measures and datasets, unlike the MLP component; classification accuracy is not always a good approximation for generation accuracy and perplexity should be calculated as well.
This study provides a comprehensive framework for creating knowledge-based hallucination datasets and offers insights into effective intervention strategies in both open-book and closed-book settings.
9 Limitations and ethics statement
This work considers a wide range of configurations for mitigating hallucinations. One limitation is that we only focused on one model (Llama2-7B) and one of its fine-tuned variants (Goat-7B). Additionally, our results span one open-book dataset (DisentQA) and one closed-book dataset (TriviaQA). Lastly, we use an initial setting and modify a single variable each time instead of checking across all options. To form a broader understanding, additional settings and models must be investigated.
The primary objective of our work is to create a better understanding of how to mitigate hallucinations. By utilizing interventions, a malicious actor could intervene in the opposite direction and increase hallucinations. However, our research aims to increase understanding and mitigation of hallucinations.
10 Acknowledgement
This research has been supported by an AI Alignment grant from Open Philanthropy, the Israel Science Foundation (grant No. 448/20), the Azrieli Foundation Early Career Faculty Fellowship, and by a grant under a Master Sponsored Research Agreement between the Technion and Google. We also thank Google Cloud for providing us with credits for running experiments on the Google Cloud Platform.
@@ A.1 General dataset construction specifics
We created at least 1,000 examples labeled as “hallucination" and 1,000 as “grounded". We filtered out instances where the golden/parametric/contextual answers exceeded 5 tokens or where the prompt exceeded 1,000 tokens.
To assess the model’s knowledge of the golden answer, we employed a beam search generation of 20 tokens, repeated 5 times with a temperature of 0.8. We utilized only examples falling under type 3 knowledge (Section 3.1), ensuring each answer was generated at least four times.
@@ A.3 Closed-book dataset construction specifics
| Prompt | Grounded | Hallucination |
@@ Appendix B Final datasets statistic
In section 3 we created a labeled dataset given a model and a dataset. DisentQA for the open-book dataset and TriviaQA for the closed-book dataset. The statistic of the final dataset count of the number of examples in each label can be found in Table 4, the else label is all the examples that did not classify to the other two labels. We can see that the different settings have different results, indicating the importance of creating a labeled dataset regarding the specific model.
| Dataset name | # Grounded | # Hallucination | # Else |
| closed-book-Llama2-TriviaQA-WACK | 4553 | 1451 | 10306 |
| open-book-Llama2-DisentQA-WACK | 1673 | 2869 | 8958 |
| closed-book-Goat-TriviaQA-WACK | 2618 | 1507 | 12185 |
| open-book-Goat-DisentQA-WACK | 3019 | 1261 | 9220 |
@@ Appendix C Specific evaluation measures
In the open book setting, we assess classification accuracy by comparing the model’s preference for the , Rank()Rank() Regarding the closed book setting, we compare Rank( | bad shots) Rank( | good shots). In both settings, the rank calculation mirrors that of the dataset creation step. If the equations are true the example is labeled ‘grounded’, else ‘hallucination’.
@@
Lastly, to calculate the Wiki ppl we use 100 random articles, and from each, we use the first 100 tokens to calculate the perplexity.
@@ Appendix D Detection results
Figures 9(a) and 9(b) show hallucination detection results on the (open-book) DisentQA-WACK and (closed-book) TriviaQA-WACK, respectively, when applied to different model components at different layers, using Llama-2-7B.
On most layers, detection is significantly higher than random (dashed lines), indicating a signal of possible hallucination before the occurrence of hallucination. Additionally, we observe that the different components yield fairly similar results, with a slight advantage for detection based on the residual stream.
We remind the reader that detecting possible hallucinations before they occur is a more challenging task than detecting them post-hallucination [5]. Indeed, if we concatenate the answer to the prompt, we reach over 90% accuracy in most layers and components, see Figure 10.101010We run this test only on the open-book setting as it has a clear answer to concatenate ( for grounded examples and for hallucination examples). Given that in practical use cases the answer is not known ahead of generation, we emphasize the importance of pre-hallucination detection.
For completeness, we present the detection results on one data split of the validation set on the residual, attention, and MLP layers in Figures 11 and 12, utilizing three different classifiers: a linear classifier trained on the vectors as they are (labeled classifier in the figures), a linear classifier trained only on the vectors’ norms (labeled norm in the figures), and a linear classifier trained on the normalized vectors (labeled direction in the figures). It’s evident that the linear classifier trained on the normalized vectors (direction classifier) achieves the best results across most components, layers, and datasets. Thus we use normalized vectors for the detection in all results.
Additionally in Figure 13 we see the classification results on the heads on Llama2-7B in the two settings. We can see that those results are a bit less good than the ones on the other model’s components with much better results in the open-book setting than the closed-book setting.
@@
So far we have performed each experiment within a given dataset and model. Do hallucination detectors generalize across datasets or models? In Figure 16(a), we employ a classifier trained on Goat-7B and evaluate its performance when detecting hallucinations from Llama2-7B states, on the DisentQA-WACK dataset (GoatLlama2; DisentQA-WACK). Interestingly, the Goat classifier consistently outperforms random chance across all layers, suggesting a notable correlation between the inner states of the two models. These results correspond to related works [43]. However, the results are much lower than a detector trained and evaluated on Llama2 states (Llama2Llama2; 9(a)), indicating that cross-model generalization is limited.
Similar to these findings, we observe a comparable outcome on the Llama2 TriviaQA-WACK setting in Figure 17. Additionally, the results of using the Llama2 model to learn a classifier for the Goat model, which is similar to the results in the other direction, also show accuracy above random, are in Figure 18. Lastly, in Figure 19, we present the results on the Goat model using each dataset to detect hallucinations on the other dataset.